OpenStudy (anonymous):

Use basic identities to simplify the expression. (1/cot^2θ) + sec θ cos θ

OpenStudy (anonymous):

Lets focus on the right side of the expression: sec0cos0 That is equal to: 1/cos0(cos0) Which is equal to: 1 So now you're left with: (1/cot^2 0) + 1 Now on the left: 1/cot^2 0 That is equal to: 1/(1/tan^20) And you can simplify that to: tan^20 Now you have: tan^20 + 1 That's equal to the identity: sec^20

OpenStudy (mathstudent55):

\(=\tan^2 \theta + \sec \theta \cos \theta\) \(=\tan^2 \theta + \dfrac{1}{\cos \theta} \cos \theta\) \(= \tan^2 \theta + 1\) \(= \sec^2 \theta\)